OpenStudy (anonymous):

Security fencing costs $5 for each 10-foot section of fence. How much would it cost to place a fence around the perimeter of a football field which measures 100 yards by 60 yards?

OpenStudy (anonymous):

perimeter of rectangle is 2L+2w = 2(100+60)= 2*160= 320yds squared convert to feet 3ft = 1 yd